# Adequately Tailoring Age Verification Regulations

### v1

@@ 6. Conclusion
In sum, age verification in the United States now sits at the intersection of child protection, privacy, free expression, and technological feasibility. The Supreme Court’s decision in FSC v. Paxton confirms that age-verification mandates are likely to persist under intermediate scrutiny; therefore laws need to be “adequately tailored” to the problem at hand (Supreme Court of the United States, 2025). However, the constitutional survival doesn’t end problems of failing to balance policy tradeoffs among governmental objectives and the need of practical implementation. To solve the remaining problems, this paper proposes a model to interpret “adequate tailoring” among government goals from measurable tradeoffs. Using the new analysis model, we evaluate current legislation and different verification implementation approaches. This approach will help with both case-by-case analysis of laws and implementations, and for drawing broader conclusions regarding trade-offs in age verification approaches.

### v2

@@ 6. Conclusion
In sum, age verification in the United States now sits at the intersection of child protection, privacy, free expression, and technological feasibility. The Supreme Court’s decision in FSC v. Paxton confirms that age-verification mandates are likely to persist under intermediate scrutiny; therefore laws need to be “adequately tailored” to the problem at hand (Supreme Court of the United States, 2025). However, the constitutional survival does not end the problem of failing to balance policy tradeoffs among governmental objectives and the need of practical implementation. To solve the remaining problems, this paper proposes a model to interpret “adequate tailoring” among government goals from measurable tradeoffs. Using the new analysis model, we evaluate current legislation and different verification implementation approaches. This approach will help with both case-by-case analysis of laws and implementations, and for drawing broader conclusions regarding trade-offs in age verification approaches.